# Dependence of the Reconstructed Core-Collapse Supernova Gravitational Wave High-Frequency Feature on the Nuclear Equation of State, in Real Interferometric Data

### v1

@@ Abstract
We present an analysis of gravitational wave (GW) predictions from five two-dimensional Core Collapse Supernova (CCSN) simulations that varied only in the Equation of State (EOS) implemented. The GW signals from these simulations are used to produce spectrograms in the absence of noise, and the emergent high-frequency feature (HFF) is found to differ quantitatively between simulations. Below 1 kHz, the HFF is well approximated by a first-order polynomial in time. The resulting slope was found to vary between 10-50% across all models. Further, using real interferometric noise we investigated the current capabilities of GW detectors to resolve these differences in HFF slope for a Galactic CCSN. We find that for distances up to 1 kpc, current detectors can resolve HFF slopes that vary by at least 30%. For further Galactic distances, current detectors are capable of distinguishing the upper and lower bounds of the HFF slope for groupings of our models that varied in EOS. With the higher sensitivity of future GW detectors, and with improved analysis of the HFF, our ability to resolve properties of the HFF will improve for all Galactic distances. This study shows the potential of using the HFF of CCSN produced GWs to provide insight into the physical processes occurring deep within CCSN during collapse, and in particular its potential to further constrain the EOS through GW detection.
first keyword, second keyword, third keyword
@@ I Introduction
From the moment of bounce onwards, each stage of a CCSN will generate GWs, as described in Abdikamalov et al. [9] and Mezzacappa and Zanolin [10]. The present study focuses on the high-frequency GW signal arising from sustained Ledoux convection due to lepton gradients deep within the PNS. These lepton gradients become large enough to cause Ledoux convection in the PNS around 100 ms postbounce and are sustained for on the order of seconds as neutrinos continue to diffuse out of the PNS. During this time, the PNS contracts. This contraction causes the frequency of the GWs produced by Ledoux convection to increase with time until the PNS reaches a stable radius, at which point the GW frequency remains constant while Ledoux convection continues. We call this feature the High Frequency Feature (HFF), using nomenclature introduced in Casallas-Lagos et al. [6]. Based on analysis of the quasi-normal modes of oscillation of the PNS, previous studies have concluded that the HFF corresponds to the frequency evolution of either a $g$-mode oscillation [15, 16, 17, 18, 19] or a combination of $g$-mode and $f$-mode oscillations [20, 21, 22, 23, 24, 25, 26]. The classification of these modes may connect to the PNS properties, as discussed in Appendix B.
@@
Early connections between GW signals in CCSNe and the EOS were noted in Seidel et al. [38]. Using an EOS that parameterized the stiffness of the EOS in spherically symetric simulations, they found that stiffer parameterizations led to slightly more efficient GW production. Marek et al. [15] investigated the GW and neutrino emissions in two-dimensional CCSN numerical simulations using a 15 $M_{\odot}$ progenitor, comparing stiff and soft EOS. They observed that PNS formed under the soft EOS, being more compact, generated GWs of higher amplitude than those produced under the stiff EOS after 100 ms post-bounce. Additionally, they noted that (i) the peak frequency of the waveform spectrum for the HFF was above 300 Hz, (ii) the peak frequency for the SASI was below 200 Hz, with (iii) both frequencies being elevated in the case of the softer EOS.
@@
Universal relations describing the GW peak frequency evolution as either an $f$- or $g$-mode were investigated in [22, 25, 41]. Bizouard et al. [42], Bruel et al. [43] investigated parameter estimation using universal relations in Gaussian simulated noise. In each study, different EOS were considered, but the focus was the accuracy of the parameterization independent of the EOS. A similar modal analysis and universal relation study for the data presented here is beyond the scope of this paper, but Appendix B shows preliminary insights into how PNS evolution may be connectable to the HFF using modal analysis.
@@
Previously, Casallas-Lagos et al. [6] formulated a machine learning (ML) approach using real LIGO interferometric noise to estimate the slope of the HFF peak over time across different two- and three-dimensional CCSN models incorporating distinct EOS. The results indicated that the HFF slope is influenced by factors such as the progenitor mass, rotation, and EOS. In scenarios with non-rotating progenitors and consistent EOS across CCSN simulations, the HFF slope without noise ranged from 3406, 1907, to 1288 Hz $\mathrm{s}^{-1}$ for progenitors of 15, 20, and 35 $M_{\odot}$, respectively, demonstrating variabilion in HFF slopes up to 63% for a mass change of 57%. For a 15 $M_{\odot}$ progenitor rotating at 0.5 rad $\mathrm{s}^{-1}$, the HFF slope found was 2246 Hz $\mathrm{s}^{-1}$, reflecting a 34% reduction from its non-rotating counterpart of the same mass. These signals were introduced into real interferometric LIGO noise through a cWB event production analysis, in standard configuration, for different Galactic distances 1, 5, and 10 kpc. The ML algorithm managed to predict the slope with an error under 10% for nearly all models except three. Moreover, the orientation of the source was found to have a minimal impact on the slope estimation of the HFF, despite variations in the intensity of CCSN GW signals with orientation, as discussed in Müller [11], Morozova et al. [21]. Differences in HFF slope due to orientation angle with the GW source were under 5%. The study confirmed the reliable estimation of the HFF slope in real interferometric noise and highlighted several physical parameters that influence this estimation.
@@
This paper continues the development of the HFF as a tool to gain physical insight from Galactic-CCSN-produced GW signals. We extract GW signals from two-dimensional, axisymmetric CCSN simulations [47, 48] performed with the Chimera  code [49]. Five models were considered that employed five distinct EOS in the CCSN simulations: DD2, FSUGold, IUFSU, SFHo, and SFHx. These EOS were chosen for their compatibility with current experimental and observational constraints, as discussed in Tews et al. [50]. We note, these EOS do not cover the full range of permissible EOS. Discussions on the explosion characteristics of the two-dimensional simulations can be found in Landfield [47]. These GW signals are first analyzed in the absence of noise to determine the intrinsic differences in HFF slope due to differing EOS in the models considered and signal processing settings. Using these intrinsic slope values, we then investigate our ability to resolve different EOS through the detection of GWs and analysis of the HFF using real interferometric LIGO data in a Coherent-Waveburst (cWB)-event production analysis [51]. Our results spans Galactic distances of 1, 5, and 10 kpc in an equatorial orientation, but varying orientations are achievable by adjusting the term $1/r$ equivalently to a multiplication by the cosine of the orientation angle $\theta$, as detailed in [52].
@@ IV HFF Detection and Analysis with Noise
Following the approach developed in [6] to predict the slope of the HFF we chose a Deep Neural Network (DNN) for our study because it is specifically designed to learn from both linear and non-linear relationships present in the input data. Additionally, it has the ability to automatically extract features from the input domain, resulting in high performance. Compared to other regression methods in the field of data analysis, the DNN offers this high performance at a relatively low computational cost. In order to assess the accuracy of the DNN algorithm, we utilized various performance metrics, which are described later in this section. Our approach for analyzing the CCSN GW signals from the E-series with interferometric noise is described as follows: (i) cWB event production: we use the cWB algorithm to create likelihood time-frequency maps, $L$, for the events that have been identified through an event production analysis. The cWB configuration parameters mirror those employed in [6], ensuring consistency in the operation of the pipeline. (ii) Likelihood Map Processing: we process the likelihood time-frequency maps to extract the physical information embedded in the detected CCSN GW events (see Section II-C in [6]). (iii) DNN Model: our primary tool for estimating the HFF slope is the DNN as described in [6].
@@ IV.3 A Deep Neural Network Model for HFF Slope Estimation in Real Interferometric Data
Table 2: HFF slope estimation results, in real interferometric noise for distances of 1 kpc, 5 kpc, and 10 kpc. For each E-series signal the slope estimation without noise is reported. Then, for each detection distance, the mean estimated slope $\hat{\bar{s}}$ is reported with its corresponding standard deviation. The root-mean-square error is given by $RMSE=\sqrt{(\sum_{i}(s_{i}-\bar{s}_{i})/n)}$, as a measure between the approximation using the slope without noise and the approximation using the mean slope from the DNN. Likewise, the MAPE uses the same formula shown in Table 1, but the error is calculated between the approximation using the slope determined without noise and the approximation using the mean slope from the DNN.
Table 2 shows the results of using the DNN algorithm for the E-series GW signals at distances of 1, 5, and 10 kpc. In each case, a mean slope $\hat{\bar{s}}$ is found using the DNN, and the RMSE and MAPE values relate the HFF approximation using $\hat{\bar{s}}$ to the value of the slope without noise $s$. Figure 9 shows the slopes of the HFF for the E-series GW signals with the ±$\pm 1\sigma$ results shaded for each model for each Galactic detection distance. Included in this figure is also a plot of the mean HFF slope $\hat{\bar{s}}$ with a range of HFF starting frequencies determined by varying $t_{\rm HFF}$ in the noiseless case.
Figure 9: The range of HFF slopes across the E-series for detection distances of 1 kpc, 5 kpc, and 10 kpc. The figures visually represent the estimation errors, which are detailed in Tables 2 for each model. The color band surrounding the data points corresponds to the STD around the mean of the estimated values $\hat{\bar{s}}$. The plot in the bottom right shows the range of HFF values centered around $\hat{\bar{s}}$, with starting frequencies varied by changing $t_{\rm HFF}$.
@@ V Discussion
Using the methods established in [6], we show the capability of using the slope of the HFF to investigate CCSNe. We note that the definition of the beginning of the HFF signal is less relevant when the signal is injected with real LIGO noise because, in this case, strains below noise levels cannot be distinguished. Thus, the HFF signal starts, in the detections with real noise, where the strain overcomes the noise threshold of the detector. Figure 9 shows the resolvability of different EOS using the slope of the HFF for 1, 5, and 10 kpc using real interferometric noise. We see that, while not all HFF slopes are distinguishable, the EOS can be divided into groups that are. The upper left plot shows that at 1 kpc, with current-generation detectors, we expect to be able to separate possible EOS into two groups, because the E-FSUGold, E-SFHo, and E-SFHx estimations remain distinguishable from the E-DD2 and E-IUFSU estimation errors. Within these groups there exist regions in the slope estimation error that would correspond to just one EOS.
Figure 9 also highlights the issue of uncertainty in the HFF starting frequency. Considering only the standard deviations in mean-slope estimates at 1 kpc in Table 2, the estimation error of the E-FSUGold HFF slope overlaps with the estimation error of E-DD2 and E-IUFSU. However, when the estimation errors of the HFF slopes are tied to a single HFF starting frequency, we see that E-FSUGold is now distinguishable from E-DD2 and E-IUFSU. Introducing the estimation error in the HFF starting frequency, in addition to the estimation error in slope, adds another parameter that will affect the resolvability of different EOS from HFF analysis, as shown in the bottom right plot in Figure 9. Even with this extra parameter space, the results are still encouraging. The comparatively large estimation error in the HFF starting frequency for E-FSUGold would cause it to overlap with the estimation error of each other model when it is combined with slope estimation error. At the same time, the HFF features of E-SFHo and E-SFHx remain distinguishable from those of E-DD2 and E-IUFSU. For Galactic distances further than 1 kpc, we leave the the estimation error in HFF starting time for future work.
@@
At 5 kpc, we can still distinguish between the same two groups of EOS based on the HFF slope but can no longer distinguish between EOS within each group. At 10 kpc, we lose the ability to distinguish between EOS based on HFF slope for all but the upper-bound slope of the E-FSUGold model and the lower-bound slope of the E-DD2 model. These results are consistent with the noise data from the LIGO O3 run conducted in 2019 through 2020. The LIGO O4 run is currently underway, with improved sensitivity, but O4 noise data is not yet available. With the improved sensitivity of future detectors, we expect the ability to measure the slope of the HFF below 1 kHz to improve as well. In our models, SNR is decreased with distance to the event due to the inverse relationship of GW signal amplitude with distance. This means that improved SNR for future detectors will effectively shift our results to larger distances by the same factor. In the case of the order-of-magnitude improvements to detection sensitivity currently projected for the Einstein Telescope [69] and Cosmic Explorer [70] detectors, our 1 kpc results would be, conservatively, representative of 10 kpc detections.

### v2

@@ Abstract
We present an analysis of gravitational wave (GW) predictions from five two-dimensional Core Collapse Supernova (CCSN) simulations that varied only in the Equation of State (EOS) implemented. The GW signals from these simulations are used to produce spectrograms in the absence of noise, and the emergent high-frequency feature (HFF) is found to differ quantitatively between simulations. Below 1 kHz, the HFF is well approximated by a first-order polynomial in time. The resulting slope was found to vary between 10–50% from model to model. Further, using real interferometric noise we investigated the current capabilities of GW detectors to resolve these differences in HFF slope for a Galactic CCSN. We find that, for distances up to 1 kpc, current detectors have the ability to resolve HFF slopes differing by 4.4–15%. For further Galactic distances, current detectors are capable of distinguishing the upper and lower bounds of the HFF slope for groupings of our models that varied in EOS. With the higher sensitivity of future GW detectors, and with improved analysis of the HFF, our ability to resolve properties of the HFF will improve for all Galactic distances. This study shows the potential of using the HFF of CCSN produced GWs to provide insight into the physical processes occurring deep within CCSN during collapse, and in particular its potential to further constrain the EOS through GW detection.
Gravitational waves, Gravitational wave detection, Novae and supernovae, Nuclear physics of explosive environments
@@ I Introduction
From the moment of bounce onwards, each stage of a CCSN will generate GWs, as described in Abdikamalov et al. [9] and Mezzacappa and Zanolin [10]. The present study focuses on the high-frequency GW signal arising from accretion onto the PNS, as well as sustained Ledoux convection due to lepton gradients deep within the PNS. These lepton gradients become large enough to cause Ledoux convection in the PNS around 100 ms postbounce and are sustained for on the order of seconds as neutrinos continue to diffuse out of the PNS. During this time, the PNS contracts. This contraction causes the frequency of the GWs produced by Ledoux convection to increase with time until the PNS reaches a stable radius, at which point the GW frequency remains constant while Ledoux convection continues. We call this feature the High Frequency Feature (HFF), using nomenclature introduced in Casallas-Lagos et al. [6]. Based on analysis of the quasi-normal modes of oscillation of the PNS, previous studies have concluded that the HFF corresponds to the frequency evolution of either a $g$-mode oscillation [15, 16, 17, 18, 19] or a combination of $g$-mode and $f$-mode oscillations [20, 21, 22, 23, 24, 25, 26, 27]. The classification of these modes may connect to the PNS properties, but such analysis is beyond the scope of this paper.
@@
Early connections between GW signals in CCSNe and the EOS were noted in Seidel et al. [39]. Using an EOS that parameterized the stiffness of the EOS in spherically symetric simulations, they found that stiffer parameterizations led to slightly more efficient GW production. Marek et al. [15] investigated the GW and neutrino emissions in two-dimensional CCSN numerical simulations using a 15 $M_{\odot}$ progenitor, comparing stiff and soft EOS. They observed that PNS formed under the soft EOS, being more compact, generated GWs of higher amplitude than those produced under the stiff EOS, after 100 ms post-bounce. Additionally, they noted that (i) the peak frequency of the waveform spectrum for the HFF was above 300 Hz, (ii) the peak frequency for the SASI was below 200 Hz, with (iii) both frequencies being elevated in the case of the softer EOS.
@@
Universal relations describing the GW peak frequency evolution as either an $f$- or $g$-mode were investigated in [23, 26, 42]. Bizouard et al. [43], Bruel et al. [44] investigated parameter estimation using universal relations in Gaussian simulated noise. In each study, different EOS were considered, but the focus was the accuracy of the parameterization independent of the EOS. A similar modal analysis and universal relation study for the data presented here could be performed but is beyond the scope of this paper.
@@
Beyond the effects of EOS, the progenitor properties can affect the HFF as well. Rotation of the progenitor tends to have a stabilizing effect on the radius of the PNS, leading to a slower contraction time. This would tend to decrease the slope of the HFF as seen, e.g., in [48, 49, 50, 51, 52]. The mass of the progenitor will affect the amplitude of the GW strains, as well as the slope of the HFF [53, 54, 55], though a systematic study is necessary to quantify the effect more precisely. Additionally, Wang and Pan [56] and Jardine et al. [57] show that the internal structure of the progenitor and the presence of strong magnetic fields can affect the HFF slope, respectively. In this study, we choose to study the effects of the EOS on the HFF in isolation of all progenitor parameters. In essence, we are assessing the maximum amount of information that could be gathered about the EOS, at least in our approach, in the event of a Galactic core collapse supernova gravitational wave detection, assuming that all of the other parameters affecting the HFF slope are known.
@@
Previously, Casallas-Lagos et al. [6] formulated a machine learning (ML) approach using real LIGO interferometric noise to estimate the slope of the HFF peak over time across different two- and three-dimensional CCSN models incorporating distinct EOS. The results indicated that the HFF slope is influenced by factors such as the progenitor mass, rotation, metalicity, and EOS. In scenarios with non-rotating progenitors and consistent EOS across CCSN simulations, the HFF slope without noise ranged from 3406, 1907, to 1288 Hz $\mathrm{s}^{-1}$ for progenitors of 15, 20, and 35 $M_{\odot}$, respectively, demonstrating variabilion in HFF slopes up to 63% for a mass change of 57%. For a 15 $M_{\odot}$ progenitor rotating at 0.5 rad $\mathrm{s}^{-1}$, the HFF slope found was 2246 Hz $\mathrm{s}^{-1}$, reflecting a 34% reduction from its non-rotating counterpart of the same mass. These signals were introduced into real interferometric LIGO noise through a cWB event production analysis, in standard configuration, for different Galactic distances 1, 5, and 10 kpc. The ML algorithm managed to predict the slope with an error under 10% for nearly all models except three. Moreover, the orientation of the source was found to have a minimal impact on the slope estimation of the HFF, despite variations in the intensity of CCSN GW signals with orientation, as discussed in Müller [11], Morozova et al. [21]. Differences in HFF slope due to orientation angle with the GW source were under 5%. The study confirmed the reliable estimation of the HFF slope in real interferometric noise and highlighted several physical parameters that influence this estimation.
@@
This paper continues the development of the HFF as a tool to gain physical insight from Galactic-CCSN-produced GW signals and, in particular, examines the effects of the EOS on it. We extract GW signals from two-dimensional, axisymmetric CCSN simulations [58, 59] performed with the Chimera  code [60]. Five models were considered that employed five distinct EOS in the CCSN simulations: DD2, FSUGold, IUFSU, SFHo, and SFHx. These EOS were chosen for their compatibility with current experimental and observational constraints, as discussed in Tews et al. [61]. We note, these EOS do not cover the full range of permissible EOS. Discussions on the explosion characteristics of the two-dimensional simulations can be found in Landfield [58]. These GW signals are first analyzed in the absence of noise to determine the intrinsic differences in HFF slope due to differing EOS in the models considered and signal processing settings. Using these intrinsic slope values, we then investigate our ability to resolve different EOS through the detection of GWs and analysis of the HFF using real interferometric LIGO data in a Coherent-Waveburst (cWB)-event production analysis [62]. Our results spans Galactic distances of 1, 5, and 10 kpc in an equatorial orientation, but varying orientations are achievable by adjusting the term $1/r$ equivalently to a multiplication by the cosine of the orientation angle $\theta$, as detailed in [63].
@@ IV HFF Detection and Analysis with Noise
Following the approach developed in [6] to predict the slope of the HFF we chose a Deep Neural Network (DNN) for our study because it is specifically designed to learn from both linear and non-linear relationships present in the input data. Additionally, it has the ability to automatically extract features from the input domain, resulting in high performance (see for example [78]). Compared to other regression methods in the field of data analysis, the DNN offers this high performance at a relatively low computational cost. In order to assess the accuracy of the DNN algorithm, we utilized various performance metrics, which are described later in this section. Our approach for analyzing the CCSN GW signals from the E-series with interferometric noise is described as follows: (i) cWB event production: we use the cWB algorithm to create likelihood time-frequency maps, $L$, for the events that have been identified through an event production analysis. The cWB configuration parameters mirror those employed in [6], ensuring consistency in the operation of the pipeline. (ii) Likelihood Map Processing: we process the likelihood time-frequency maps to extract the physical information embedded in the detected CCSN GW events (see Section II-C in [6]). (iii) DNN Model: our primary tool for estimating the HFF slope is the DNN as described in [6]. For a comparison of different attempts to perform parameter estimation for core-collapse supernovae we refer the reader to [79, 80].
@@ IV.3 A Deep Neural Network Model for HFF Slope Estimation in Real Interferometric Data
Table 2: HFF slope estimation results in real interferometric noise for distances of 1 kpc, 5 kpc, and 10 kpc. The slope estimation without noise is reported for each E-series signal in the second column. Then, for each detection distance, the mean estimated slope $\hat{\bar{s}}$ is reported with its corresponding standard deviation. The root-mean-square error is given by $RMSE=\sqrt{(\sum_{i}(s_{i}-\bar{s}_{i})/n)}$, as a measure between the approximation using the slope without noise and the approximation using the mean slope from the DNN. Likewise, the MAPE uses the same formula shown in Table 1, but the error is calculated between the approximation using the slope determined without noise and the approximation using the mean slope from the DNN.
Table 2 shows the results of using the DNN algorithm for the E-series GW signals at distances of 1, 5, and 10 kpc. In each case, a mean slope $\hat{\bar{s}}$ is found using the DNN, and the RMSE and MAPE values relate the HFF approximation using $\hat{\bar{s}}$ to the value of the slope without noise $s$. Table 3 reports the same results for the HFF starting frequency, denoted as $\beta$ in Section III.1.
1 kpc 5 kpc 10 kpc
EOS $\bar{\beta}$ STD RMSE MAPE $\bar{\beta}$ STD RMSE MAPE $\bar{\beta}$ STD RMSE MAPE
[Hz] [Hz] [Hz] [%] [Hz] [Hz] [Hz] [%] [Hz] [Hz] [Hz] [%]
DD2 419 5.21 11.23 4.32 531 12.94 13.54 10.18 649 27.39 31.30 23.62
FSUGold 429 8.51 13.47 6.05 550 18.92 27.14 15.73 714 31.32 34.01 30.98
IUFSU 378 9.01 14.04 6.67 505 20.04 29.37 16.34 708 33.51 36.37 33.75
SFHo 364 8.39 12.57 6.01 481 19.49 27.01 15.24 698 30.21 33.23 33.13
SFHx 408 5.87 11.74 4.94 523 15.50 25.21 11.36 660 29.94 32.12 26.17
Table 3: HFF starting frequency estimation results, in real interferometric noise, for distances of 1 kpc, 5 kpc, and 10 kpc. $\bar{\beta}$ is the mean HFF starting frequency obtained from the DNN, and it is reported with its corresponding standard deviation for each Galactic distance. The RMSE and MAPE are computed using the same formulas described in Table 2 and Table 1, respectively. These quantities are computed with respect to their values as determined without noise.
Figure 9 shows the linear estimation of the HFF for the E-series GW signals with the ±$\pm 1\sigma$ results shaded for each model for each Galactic detection distance. This shading encompasses the standard deviations in both HFF slope and HFF starting frequency. As seen in Table 3, the mean HFF starting frequency varies with detection distance. This is reflected in Figure 9, as each plot has a different HFF starting frequency on the frequency axis.
Figure 9: HFF slopes estimated by our DNN model including the range of variability of the HFF starting frequency across E-series models for detection distances of 1 kpc, 5 kpc, and 10 kpc. The figures visually represent the estimation errors in both HFF slope and starting frequency. The color band surrounding the data points corresponds to the STD around the mean of the estimated values of $\hat{\bar{s}}$ and $\bar{\beta}$ as is reported in table 2 and 3.
@@ V Discussion
Using the methods established in [6], we show the capability of using the slope of the HFF to investigate CCSNe. We note that the definition of the beginning of the HFF signal is less relevant when the signal is injected with real LIGO noise because, in this case, strains below noise levels cannot be distinguished. Thus, the HFF signal starts, in the detections with real noise, where the strain overcomes the noise threshold of the detector. The mean HFF starting frequency reported in Table 3 is then a result of the range of starting frequencies that may be detected depending on when the signal overcomes the noise threshold.
Figure 9 shows the estimation errors for both HFF slope and HFF starting frequency for 1, 5, and 10 kpc. At 1 kpc, there are regions with no overlap in estimation error between each E-series model. We see the emergence of two groupings: the steep-sloped E-SFHo and E-SFHx models and the more-shallow-sloped E-DD2 and E-IUFSU models. The estimation error of E-FSUGold crosses both groupings, having a high HFF starting frequency but more shallow slope. Having low estimation error in the HFF starting frequency allows for clear distinctions between models with similar slopes. E-SFHo and SFHx are similar in slope, but well resolved and separated in HFF starting frequency. The same is true for E-IUFSU and E-DD2. The minimum slope resolvability is between E-SFHo and E-SFHx, which differ by 4.4% and have HFF starting frequencies differing by 12%. The minimum HFF starting frequency resolvability is between E-DD2 and E-FSUGold, which differ by 2.4% and have HFF slopes differing by 15%. In other words, for EOS models with larger differences in HFF starting frequency we can resolve smaller differences in HFF slope. Conversely, for EOS models with roughly similar HFF starting frequencies we require a greater separation in HFF slope in order to resolve them.
At 5 kpc, the HFF starting frequency is notably higher across all models. This is due to the weaker signal strength at the beginning of the HFF when compared to the increased noise, so the HFF is detected at a later time in the signal. We thus lose the clear period of resolvability at early signal times, as seen in the 1 kpc detection. Nonetheless, E-SFHx and E-SFHo remain largely distinguishable from E-DD2 and E-IUFSU, though distinguishing within these groups is more difficult. E-FSUGold remains distinguishable at early times, but not at later times.
At 10 kpc, as the the reconstruction of the earlier HFF signal is further affected by the noise, the HFF starting frequency continues to rise. The resolvability between the groups, seen at 1 and 5 kpc, is lost. However, the upper bounds of E-FSUGOld and lower bounds of E-DD2 in both HFF starting frequency and slope are still resolvable. With the expected improvements in sensitivity in future runs, when interferometers will reach design sensitivity and when the number of interferometers will increase, we expect the uncertainty in the starting frequency to decrease, as well as the uncertainty in the slope. For this reason the results presented here should be thought of as conservative.
@@
These results are consistent with the noise data from the LIGO O3 run conducted in 2019 through 2020. The LIGO O4 run is currently underway, with improved sensitivity, but O4 noise data is not yet available. With the improved sensitivity of future detectors, we expect the ability to measure the slope of the HFF below 1 kHz to improve as well. In our models, SNR is decreased with distance to the event due to the inverse relationship of GW signal amplitude with distance. This means that improved SNR for future detectors will effectively shift our results to larger distances by the same factor. In the case of the order-of-magnitude improvements to detection sensitivity currently projected for the Einstein Telescope [82] and Cosmic Explorer [83], our 1 kpc results would be, conservatively, representative of 10 kpc detections.